A person driving a car with a speed 72 km/hr, suddenly sees a boy crossing the road. If the distance moved by car, before the person applies brakes is 5 m, the reaction time of the person is1) 0.5 s2)0.15 s3) 10 s4) 0.25 sBy using equations of kinetmaics I am getting option 1 as answer but by applying distance = speed*time I am getting option 4 as answer.Option 4 is the answer but why we wont use equations of kinematics explain please?

Grade:12th pass

3 Answers

Amritpal Singh Saini
5 years ago
Dear student, Kinematics equation are often used in a condition when the motion is described by a constant velocity (acceleration is 0) , they cannot be used on a situation where in the acceleration is changing! Whereas when the object moves in the Straight line with steady speed , u can use speed distance relation. Provided that u have the necessary information! (Please do approve my answer)
Kervender kaur oberoi
5 years ago
U= 20 m/s ...S= 5 m....So..
20 m dist.covered in 1sec. 
1m dist.covered in 1/20 sec.
5.m dist.covered in 1/20×5 ...= 0.25 sec.
Vanshika
5 years ago
72 kilometer per hour = 20 metre per second
Distance = Velocity * Time
5= 20*Time
T= 5/20
   =1/4
   =0.25 seconds